OpenStudy (chupacabraj):

I'm having a little trouble finding the critical points for f(x,y) = (x-y)(1-xy)

OpenStudy (chupacabraj):

I expanded it---> f(x,y) = x-x^2x - y +xy^2

OpenStudy (chupacabraj):

\[f_x = 1-2xy+y^2 = 0\]

OpenStudy (chupacabraj):

\[f_y= -x^2 -1+2xy=0\]

OpenStudy (chupacabraj):

I'm not sure how to go about it. I attempted to solve using elimination and got x^2=y^2

OpenStudy (seratul):

I don't really know how to solve this, but you made a mistake with your math. It should be f(x,y) = x-x^2y-y xy^2

OpenStudy (seratul):

f(x,y) = x-x^2y-y+xy^2

OpenStudy (chupacabraj):

I got x^2=y^2

OpenStudy (triciaal):